A ball is dropped from a height of 1m. How long does it take for the ball to reach theground? With what speed does it hit the ground?


I GOT ANSWERS=0.14s;1.4 m/s respectively.


IS not my answer correst? BOOK HAS ANSWER PRINTED 0.45s; 4.4m/s.

Grade:10

3 Answers

hari krishn
11 years ago

initial velocity is zero so final velocity when it reaches ground is sqrt(2gh), where h is the height of fall......so velocity sqrt(2*9.8*1)=4.4 m/s

 

the time taken for fall t=velcity divided by acceleration........t=4.4/9.8=0.45 seconds.........

shritesh srivastava
11 years ago

dear friend,the book is correct.                                                                                                                                        since u=0,h=1m    .now apply 2nd and 3rd law..you will get the correct answer.  

suru siva
11 years ago

v^2=2*g

v^2=20

v=4.4m/s

t=v/g

t=.44s